OpenStudy (anonymous):

Sand is being poured at a rate of 2 m^3/second onto a conical pile of sand that is initially 2 meters in height with a base diameter of 6 meters. How fast is the height of the pile increasing at the moment when the pile of sand is 4 meters in height?

OpenStudy (anonymous):

dv/dt= 2 dh/dt=? v= 1/3 Pi r^2 h dv= 1/3 Pi r^2 dh

OpenStudy (anonymous):

2=1/3 Pi (3)^2 dh use algebra to find dh